# Demonstration of Efficient Radon Removal by Silver-Zeolite in a Dark Matter Detector

### v1

@@ Abstract
We present the performance of an efficient radon trap using silver-zeolite Ag-ETS-10, measured with a spherical proportional counter filled with an argon/methane mixture. Our study compares the radon reduction capabilities of silver-zeolite and the widely used activated charcoal, both at room temperature. We demonstrate that silver-zeolite significantly outperforms activated charcoal by three orders of magnitude in radon capture. Given that radon is a major background contaminant in rare event searches, our findings highlight silver-zeolite as a highly promising adsorbent, offering compelling operational advantages for both current and future dark matter and neutrino physics experiments. Additionally, this holds great promise for developing future radon reduction systems in underground laboratories in order to increase their supporting capabilities for ton-scale experiments.
@@
Radon ( [^2] Rn), a naturally occurring radioactive noble gas produced from the decay of 238U, poses significant challenges to low-background detection techniques in particle physics. As a major contaminant, radon compromises the sensitivity of experiments such as dark matter searches and neutrino physics studies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39].  [^2] Rn decays by emitting an alpha particle with an energy of 5.49 MeV and has a half-life of 3.82 days [40]. Its decay chain includes two other alpha emitters, 218Po and 214Po, before finally producing 210Pb, which is also a significant background source [41, 42, 43, 44, 19, 31]. As a noble gas, radon can continuously emanate from the detector material and other components such as gas filters and pumps, and subsequently diffuse through the active target. Beta emitters from the  [^2] Rn decay chain, such as 214Pb and 210Pb, can also contribute to a homogeneous background of low-energy events in the region of interest for rare event searches that cannot be discriminated against in data analysis [41, 42, 43, 45]. Although existing cryogenic methods—such as activated charcoal or copper traps [1, 46, 16, 47, 48, 49, 50] and distillation [51, 45, 35]—have shown good radon removal performance, the increasing complexity and expansion to ton-scale experimental setups [52, 53, 54, 55, 18, 56, 57, 58, 59, 60, 61, 62, 63, 37, 38, 64, 65, 66] require the development of more efficient and practically optimized removal techniques.
This Letter presents a radon trap system developed for the New Experiments with Spheres-Gas (NEWS-G) dark matter experiment, currently achieving world-leading dark matter sensitivity [67, 68]. The detector consists of a spherical proportional counter (SPC) filled with a noble gas mixture. It is equipped with a high-voltage (HV) sensor in the middle to achieve high avalanche amplification gain while having a linear energy response and low electronic noise [69, 70]. As a gaseous ionization detector, an SPC is capable of observing the energy response of single electrons [71], which makes it an ideal detector for searching for sub-GeV/$c^{2}$ dark matter candidates. However, electronegative impurities in the gas—such as water and oxygen—absorb primary electrons and compromise energy signal reconstruction through a process called attachment [72, 73]. To address this, the NEWS-G gas handling system includes a gas purifier (or getter) to remove these impurities. Unfortunately, the getter also releases significant amounts of radon, with its daughter elements depositing on the inner surface of the SPC and causing low-energy background events in the region of interest for dark matter searches [67]. Traditional methods like active detector shielding and fiducialization cannot fully reject them.
@@
For our study, we evaluate radon removal performance of silver-zeolite, specifically Ag-ETS-10 [74, 75], compared to the commonly used coconut shell granular activated charcoal (Silcarbon), both tested at room temperature. The rate of radon and its daughter emitters are monitored using a small-scale SPC detector filled with an argon/methane mixture, integrated within a closed-loop circulation system. We show that, at room temperature, the silver-zeolite significantly outperforms activated charcoal by three orders of magnitude in radon capture, with near-complete removal of radon from the system.
@@
Our work considerably extends previous studies on radon trapping with silver-zeolite Ag-ETS-10. The efficacy of various silver-exchanged zeolites has been demonstrated at room temperature in the important works [76, 77, 78, 79, 80, 81], with silver-zeolite Ag-ETS-10 consistently emerging as one of the most promising adsorbents for radon mitigation [77, 78]. These earlier studies on Ag-ETS-10 have been conducted with open-loop systems using gases such as nitrogen, air, or argon and employed a variety of detection methods, including the RadonEye detector and the alphaGUARD monitor. While those detectors are efficient for radon measurement, our work constitutes the first in-situ application of a silver-zeolite trap implemented in a dark matter direct detection-type experiment using an SPC. SPCs offer better energy resolution and sensitivity [82, 69, 83], with the ability to present our findings near background-free in contrast to other methods. Our unique experimental setup shows how the performance of silver-zeolite Ag-ETS-10 traps for radon removal would be in dark matter detectors like NEWS-G and other rare event searches experiments, providing a distinct advantage over other configurations.
@@
Figure 1: Closed-loop circulation system for the radon removal campaigns, made of stainless steel fittings, valves and tubing. The setup includes a radon source to diffuse radon into the SPC detector. The gas from the detector (red) circulates through the trap, where radon is adsorbed. The purified gas (blue) is pumped back to the detector, passing through a particle filter, a binary gas analyzer (BGA), and a custom-made laser absorption spectroscopy system (LAS) [76, 84].
@@
Experimental setup—We monitor the event detection rate in a 30 cm diameter stainless steel SPC filled with 500 mbar of a 97% argon and 3% methane mixture. The detector is equipped with a two-mm diameter spherical anode at its center, biased to a positive HV of 1180 V via an insulated HV wire going through the grounded support rod. The operating conditions are selected to contain the track lengths of alpha particles (210Po (8.62 cm),  [^2] Rn (9.09 cm), 218Po (10.3 cm), 214Po (15.1 cm), calculated with SRIM [85]) so that these deposit their full kinetic energy in the gas volume with high amplification gain. During each measurement, the SPC is integrated into a closed-loop circulation system consisting of a radon source, a radon trap, and a pump with a flow of 1 L/min, as shown in Fig. 1. Continuous circulation ensures the retention of radon in the trap until disintegration, while the purified gas is recirculated back into the detector.
@@
The radon source (Pylon 1025) uses dry 226Ra to produce a calibrated quantity of  [^2] Rn, with an equilibrium activity of 0.93 kBq and a continuous rate of 117.49 mBq. The radon trap consists of a custom 20 cm long stainless steel 1/2-inch tube, filled with 10 g of adsorbent material (silver-zeolite or activated charcoal) and capped at both ends with multi-purpose fiberglass. In addition, a particle filter (60 ${\upmu}$m pore size) is installed after the trap to prevent any granules of the adsorbent material from entering sensitive equipment within the circulation system. A binary gas analyzer (BGA) and a custom-made laser absorption spectroscopy system (LAS) are included to monitor the methane concentration in the gas during measurements [76, 84].
@@
Before each measurement, the trap undergoes thermal regeneration at 160C within the closed-loop system. This process involves flushing with nitrogen gas via the gas inlet to the exhaust, followed by evacuation with a vacuum pump and isolation of the trap [86]. This regeneration, also known as activation, removes any adsorbed gases and contaminants. To further minimize contaminants in the gas loop, the entire system is evacuated to approximately $10^{-7}$ Torr, except the circulation pump, which cannot achieve a vacuum lower than $10^{-3}$ Torr. Finally, the gas mixture is injected, filling the system to a pressure of 500 mbar.
@@
Data analysis—To interpret the data and assess the experimental results, we apply one of two distinct cuts to the data. The first is a quality cut to specifically remove events from atmospheric muons, which deposit tens of keV in the detector, as well as electronic noise events. The result of applying this cut in all phases of the three campaigns is shown in Fig. 2. The constant background observed in phase I comes predominantly from the alpha decay of 210Po (5.3 MeV [87]), originating from the stainless-steel inner surface of the SPC detector. It serves as a stable and constant low-rate calibration source for this analysis [86].
The second is the 214Po cut to select only 214Po events. When the radon is injected in phases II and III, the amplitude distribution exhibits different peaks corresponding to  [^2] Rn (5.5 MeV [40]) and the daughter isotopes 218Po (6.0 MeV [88]) and 214Po (7.7 MeV [89]), proportional to the energy of each alpha emitter (see Fig. S1 of [86]). This series of amplitude peaks is used to define a more strict selection of only full energy deposition 214Po events, based on the amplitude minima between the 214Po and 218Po peaks. However, since no 214Po peak features appear in phase I for all campaigns and in phase IV for campaigns 1 and 2, the 214Po cuts in these phases are derived from the 210Po peak and scaled assuming a linear energy response [86]. As shown in Fig. 2, the selection of only 214Po events significantly improves our signal-to-background ratio and allows us to assess our results as near background-free in later analysis.
The expected  [^2] Rn decay rate over time without the intervention of the trap, with either the quality or 214Po cut applied, is determined using the phase III data of each campaign. High event rates ($\sim$ 100 Hz) in the SPC cause pileup and dead-time losses of alpha events [90]. To infer the true radon rate from the measured rate in-spite of these factors, a Monte Carlo (MC) simulation is designed to emulate the decay process and include expected pileup and dead-time effects [86]. This model is fit to the data using a Markov Chain Monte Carlo (MCMC) [91], jointly incorporating the observed rate in phase III with either the quality or 214Po cut applied. The resulting fit of the observed event rates with each cut applied is shown in Fig. 2, including extrapolations to phase IV for each campaign to show the expected decay rates without the intervention of the radon trap.
@@
Rate comparison—Figure 2 clearly demonstrates that campaigns 1 and 2, which utilize the silver-zeolite trap, outperform campaign 3 using activated charcoal in terms of radon and 214Po event reduction. In particular, in campaigns 1 and 2, the rates in phase IV drop significantly before stabilizing. In contrast, the rate remains higher in phase IV of campaign 3, with an exponential decay indicating the presence of radon remaining in the detector.
As an initial assessment of the trap performance, we compare the rates obtained in phases I and IV with either quality or 214Po cut throughout the three campaigns. The rate in phase I for all campaigns is approximately constant, and is fitted with a constant function. To compare this rate with the rate in phase IV after the trap is opened, we define the following reference time. We extrapolate the expected rate from phase III (without trap) with quality cut to later times and determine the reference time where the extrapolation reaches 76.5 Hz (dashed vertical lines in Fig. 2). This procedure is applied in all three campaigns to determine the rates in phases I and IV, see Table S1 of [86]. From this, we calculate the rate ratio between phase IV and phase I with either the quality or 214Po cut, as shown in Fig. 3.
Campaigns 1 and 2, using the silver-zeolite trap, consistently achieve near complete radon removal, whereas the radon levels in the activated charcoal campaign 3 are two to three orders of magnitude higher. As expected, the rate ratio obtained with the 214Po cut is typically larger, but more affected by statistical fluctuations due to the small rate of events in the background measurement of phase I.
@@
Here, $R_{\text{exp}}(t)$ is the expected rate obtained from the extrapolated rate trend and $R_{\text{obs}}(t)$ is the observed rate from the experimental data after the trap is opened. The R-value for each time bin in phase IV is determined by optimizing (with a Nelder-Mead algorithm [92, 93]) a Poisson likelihood comparing the observed count rate with the expected count rate scaled by $R(t)^{-1}$ [86]. The Feldman-Cousins method is applied to determine the 90% confidence level (CL) uncertainty bands for $R(t)^{-1}$ [94]. A rolling-average smoothing over 15 time bins is applied to the best-fit $R(t)^{-1}$ and the 90% CL intervals to mitigate statistical fluctuations in the post-trap data. This process is repeated with many MCMC samples for the extrapolated rate trend model, and the resulting median value of the best-fit $R(t)^{-1}$ and the 90% CL interval is taken in each time bin.
Given the average track length of 214Po events (15.1 cm in the relevant gas conditions [85]) relative to the 30 cm diameter of the SPC, there is the possibility that some events may not deposit all their energy within the detector volume. Further, some fraction of 214Po isotopes are expected to settle on the inner surface of the detector before decaying, increasing the fraction of events that will not deposit their full energy in the SPC gas. This could result in an underestimation of the actual rate of 214Po. Consequently, the expected and observed 214Po event rates in phase IV are both conservatively reduced by 50% in the calculation of $R(t)^{-1}$.
To quantitatively compare the R-value across all campaigns, we extract the 90% lower confidence limit (LCL) of $R$ at the reference time (see Fig. S3 in [86]). The results are presented in Fig. 4 with the quality cut or the 214Po cut (with 50% reduced signal efficiency) applied. The corresponding values are reported in Table S2 of [86]. In campaigns utilizing silver-zeolite (campaigns 1 and 2), the 90% LCL R-values range from 3.2$\times$103 to 6.4$\times$103, which is three orders of magnitude higher than those found in campaign 3, using activated charcoal, where R-values range from 5.3 to 11.6. Moreover, campaigns 1 and 2 demonstrate consistent performance of silver-zeolite, as evidenced by their similar results. As anticipated, R-values derived with the 214Po cut applied are generally higher than those with the quality cut, due to a higher signal-to-background ratio.
@@
In this Letter, we demonstrated that silver-zeolite Ag-ETS-10 exhibits superior radon removal efficiency at room temperature compared to activated charcoal, which is commonly used at cryogenic temperatures in rare event search experiments. Our findings highlight the potential of silver-zeolite for developing ambient-temperature radon reduction systems. This advancement is particularly important for underground laboratories, where elevated radon levels compromise the ultra-low background environments required for dark matter and neutrino physics experiments [95, 96]. Integrating silver-zeolite filters into ventilation systems and clean rooms can substantially reduce radon concentrations and enhance the supporting capabilities of these facilities [97, 98]. Moreover, eliminating the need for cryogenic cooling offers significant operational advantages by simplifying system design and reducing the overall complexity of radon reduction systems for future ton-scale detectors. Since our experimental campaigns with silver-zeolite Ag-ETS-10 have been conducted with an actual dark matter-type detector, they serve as a proof of principle for its application in current and future rare event searches.
Acknowledgments—This work was supported by the Canada First Research Excellence Fund through the Arthur B. McDonald Canadian Astroparticle Physics Research Institute and through the University of Alberta startup fund UOFAB Startup Piro. M.-C. P. also acknowledges funding from the Natural Sciences and Engineering Research Council of Canada (NSERC) Subatomic Physics Discovery Grant (project) No. SAPPJ-2024-00025, from which P. G. and Y. D. were further supported. The work of D. D. and C. G. was also supported by the NSERC Graduate Scholarships Doctoral and Master programs, respectively. We are grateful for the support of Compute Ontario [99], the BC DRI Group, and the Digital Research Alliance of Canada [100] required to carry out the simulation needed for this Letter. We thank Hervé Le Provost (IRFU-CEA/Saclay) for his software development work in our data acquisition system. We also thank the NEWS-G collaboration for their technical support of the experimental setup provided for this work and the review of this Letter. We also gratefully thank Daniel Kuznicki for assisting with the silver-zeolite Ag-ETS-10 sample obtained from Extraordinary Adsorbents Inc. of Edmonton, Alberta [74].
@@
This Supplemental Material provides further technical details regarding the regeneration process of the absorbents, the cuts used in our data analysis, the expected decay rates derived with a Monte Carlo (MC) model and Markov Chain Monte Carlo (MCMC) fit, the rates in phases I and IV obtained throughout the campaigns, and calculation of the R-values.

### v2

@@ Abstract
We present the performance of an efficient radon trap using silver-zeolite Ag-ETS-10, measured with a spherical proportional counter filled with an argon/methane mixture. Our study compares the radon reduction capabilities of silver-zeolite and the widely used activated charcoal, both at room temperature. We demonstrate that silver-zeolite significantly outperforms activated charcoal by three orders of magnitude in radon capture. Given that radon is a major background contaminant in rare event searches, our findings highlight silver-zeolite as a highly promising adsorbent, offering compelling operational advantages for both current and future dark matter and neutrino physics experiments. Furthermore, this not only offers great promise for developing future radon reduction systems in underground laboratories, but also paves the way for innovative, multidisciplinary advancements with far‑reaching implications in science, engineering and environmental health.
@@
Radon ( [^2] Rn), a naturally occurring radioactive noble gas produced from the decay of 238U, poses significant challenges to low-background detection techniques in particle physics. As a major contaminant, radon compromises the sensitivity of experiments such as dark matter searches and neutrino physics studies [ODwyer:2011, Pollmann:2012, GERDA:2013krg, Amaudruz:2012hr, Nakano:2017rsy, Miller:2017tpl, Benato:2017kdf, Rupp:2017zcy, Street:2017bde, Quintana:2018, Monte:2018vjt, NEXT:2018zho, CRESST:2019oqe, Lehnert:2018cpv, Zhang:2020qfs, LZ:2020fty, Borexino:2019wln, nEXO:2021ujk, DAMIC:2021crr, XENON:2021mrg, Hussain:2022, Arthurs:2022, PandaX-4T:2021lbm, Ha:2022psk, SNO:2022qvw, Street:2023, CUPID-Mo:2023vle, LZ:2022ysc, LZ:2022lsv, Chott:2022lnc, Veeraraghavan:2023vxd, MicroBooNE:2023sxs, Seo:2024rea, XENON:2024lbh, Cui:2024ltd, MillerChikowski:2024, XLZD:2024nsu, XLZD:2024pdv, Lahaie:2024jty].  [^2] Rn decays by emitting an alpha particle with an energy of 5.49 MeV and has a half-life of 3.82 days [nndc_218]. Its decay chain includes two other alpha emitters, 218Po and 214Po, before finally producing 210Pb, which is also a significant background source [XMASS:2017sid, Bunker:2020sxw, NEWS-G:2020fhm, Vivo-Vilches:2021lwr, DAMIC:2021crr, Veeraraghavan:2023vxd]. As a noble gas, radon can continuously emanate from the detector material and other components such as gas filters and pumps, and subsequently diffuse through the active target. Beta emitters from the  [^2] Rn decay chain, such as 214Pb and 210Pb, can also contribute to a homogeneous background of low-energy events in the region of interest for rare event searches, posing significant challenges to discriminate against [XMASS:2017sid, Bunker:2020sxw, NEWS-G:2020fhm]; however, data analysis approaches have shown promising success in mitigating their impact [XENON:2024lbh, LZ:2024zvo]. Although existing cryogenic methods—such as activated charcoal or copper traps [ODwyer:2011, ABE201250, LZ:2020fty, Chen:2021hva, MicroBooNE:2022his, Li:2023lvs, Fatemighomi:2025gqd] and distillation [XENON100:2017gsw, Murra:2022mlr, Cui:2024ltd, XENON:2025nic]—have shown good radon removal performance, the increasing complexity and expansion to ton-scale experimental setups [DarkSide-20k:2017zyg, Vahsen:2020pzb, Garcia-Viltres:2021swf, SNO:2021xpa, nEXO:2021ujk, JUNO:2021vlw, CUPID:2022wpt, KamLAND-Zen:2022tow, NEWS-G:2023qwh, Novella:2023gqb, Abe:2024kjq, McDonald:2024osu, GlobalArgonDarkMatter:2024wtv, XLZD:2024nsu, XLZD:2024pdv, PANDA-X:2024dlo, DUNE:2024wvj, Calgaro:2024boi, Bouet:2024neu] require the development of more efficient and practically optimized removal techniques.
[…]
@@
For our study, we evaluate radon removal performance of silver-zeolite, specifically Ag-ETS-10 [ExtraEdmonton1, ExtraEdmonton2], compared to the commonly used coconut shell granular activated charcoal (Silcarbon), both tested at room temperature. The rate of radon and its daughter emitters are monitored using a small-scale SPC detector filled with an argon/methane mixture, integrated within a closed-loop circulation system. We show that, at room temperature, the silver-zeolite significantly outperforms activated charcoal by three orders of magnitude in radon capture, with near-complete removal of radon from the system.
@@
Our work considerably extends previous studies on radon trapping with silver-zeolite Ag-ETS-10. The efficacy of various silver-exchanged zeolites has been demonstrated at room temperature in the important works [Obrien:2021, Heinitz:2023, 10.1093/ptep/ptad160, Ogawa:2022tku, Sone:2024gox, 10.1093/ptep/ptaf007], with silver-zeolite Ag-ETS-10 consistently emerging as one of the most promising adsorbents for radon mitigation [Heinitz:2023, 10.1093/ptep/ptad160]. These earlier studies on Ag-ETS-10 have been conducted with open-loop systems using gases such as nitrogen, air, or argon and employed a variety of detection methods, including the RadonEye detector and the alphaGUARD monitor. While those detectors are efficient for radon measurement, our work constitutes the first in-situ application of a silver-zeolite trap to validate its performance under realistic operating conditions using in a dark matter direct detection-type experiment with an SPC. SPCs offer better energy resolution and sensitivity [Giomataris:2008ap, Brossard:2019eby, Giomataris:2022bvz], with the ability to present our findings near background-free in contrast to other methods. Moreover, our novel closed-loop experimental setup, in contrast to previous methods, shows how the performance of silver-zeolite Ag-ETS-10 traps for radon removal would be in dark matter detectors like NEWS-G and other rare event searches experiments, providing a distinct advantage over other configurations.
@@
Figure 1: Closed-loop circulation system for the radon removal campaigns, made of stainless steel fittings, valves and tubing. The setup includes a radon source to diffuse radon into the SPC detector. The gas from the detector (red) circulates through the trap, where radon is adsorbed. The purified gas (blue) is pumped back to the detector, passing through a particle filter, a binary gas analyzer (BGA), and a custom-made laser absorption spectroscopy system (LAS) [Obrien:2021, Garrah:2023].
@@
Experimental setup—We monitor the event detection rate in a 30 cm diameter stainless steel SPC filled with 500 mbar of a 97% argon and 3% methane mixture. The detector is equipped with a two-mm diameter spherical anode at its center, biased to a positive HV of 1180 V via an insulated HV wire going through the grounded support rod. The operating conditions are selected to mostly contain the track lengths of alpha particles (210Po (8.62 cm),  [^2] Rn (9.09 cm), 218Po (10.3 cm), 214Po (15.1 cm), calculated with SRIM [Ziegler:2010]) so that these deposit their full kinetic energy in the gas volume with high amplification gain. During each measurement, the SPC is integrated into a closed-loop circulation system consisting of a radon source, a radon trap, and a pump with a flow of 1 L/min, as shown in Fig. 1. The circulation ensures sustained radon capture by continuously passing the gas through the trap, while the purified gas is subsequently recirculated back into the detector.
@@
The radon source (Pylon 1025) uses dry 226Ra to produce a calibrated quantity of  [^2] Rn, with an equilibrium activity of 0.93 kBq and a continuous rate of 117.49 mBq. The radon trap consists of a custom 20 cm long stainless steel 1/2-inch tube, filled with 10 g of adsorbent material (silver-zeolite or activated charcoal) and capped at both ends with multi-purpose fiberglass. In addition, a particle filter (60 $\upmu$}m pore size) is installed after the trap to prevent any granules of the adsorbent material from entering sensitive equipment within the circulation system. A binary gas analyzer (BGA) and a custom-made laser absorption spectroscopy system (LAS) are included to monitor the methane concentration in the gas during measurements [Obrien:2021, Garrah:2023].
@@
Before each measurement, the trap undergoes thermal regeneration at 160C within the closed-loop system. This process involves flushing with nitrogen gas via the gas inlet to the exhaust, followed by evacuation with a vacuum pump and isolation of the trap [SOM]. This regeneration, also known as activation, removes any adsorbed gases and contaminants. To further minimize contaminants in the gas loop, the entire system is evacuated to approximately $10^{-7}$ Torr, except the circulation pump, which cannot achieve a vacuum lower than $10^{-3}$ Torr. Finally, the gas mixture is injected, filling the system to a pressure of 500 mbar.
@@
Data analysis—To interpret the data and assess the experimental results, we apply one of two distinct cuts to the data. The first is a quality cut to specifically remove events from atmospheric muons, which deposit tens of keV in the detector, as well as electronic noise events. The result of applying this cut in all phases of the three campaigns is shown in Fig. 2. The constant background observed in phase I comes predominantly from the alpha decay of 210Po (5.3 MeV [nndc_206]), originating from the stainless-steel inner surface of the SPC detector. It serves as a stable and constant low-rate calibration source for this analysis [SOM].
The second is the 214Po cut to select only 214Po events. When the radon is injected in phases II and III, the amplitude distribution exhibits different peaks corresponding to  [^2] Rn (5.5 MeV [nndc_218]) and the daughter isotopes 218Po (6.0 MeV [nndc_214]) and 214Po (7.7 MeV [nndc_210]), proportional to the energy of each alpha emitter (see Fig. S1 of [SOM]). This series of amplitude peaks is used to define a more strict selection of only full-energy-deposition 214Po events, based on the amplitude minima between the 214Po and 218Po peaks. However, since no 214Po peak features appear in phase I for all campaigns and in phase IV for campaigns 1 and 2, the 214Po cuts in these phases are derived from the 210Po peak and scaled assuming a linear energy response [SOM]. As shown in Fig. 2, the selection of only 214Po events significantly improves our signal-to-background ratio and allows us to assess our results with near-zero background in later analyses.
The expected  [^2] Rn decay rate over time without the intervention of the trap, with either the quality or 214Po cut applied, is determined using the phase III data of each campaign. High event rates ($\sim$ 100 Hz) in the SPC cause pileup and dead-time losses of alpha events [Durnford:2024]. To infer the true radon rate from the measured rate in-spite of these factors, a Monte Carlo (MC) simulation is designed to emulate the decay process and include expected pileup and dead-time effects [SOM]. This model is fit to the data using a Markov Chain Monte Carlo (MCMC) [Foreman-Mackey:2012any], jointly incorporating the observed rate in phase III with either the quality or 214Po cut applied. The resulting fit of the observed event rates with each cut applied is shown in Fig. 2, including extrapolations to phase IV for each campaign to show the expected decay rates without the intervention of the radon trap.
@@
Rate comparison—Figure 2 clearly demonstrates that campaigns 1 and 2, which utilize the silver-zeolite trap, outperform campaign 3 using activated charcoal in terms of radon and 214Po event reduction. In particular, in campaigns 1 and 2, the rates in phase IV drop significantly before stabilizing. In contrast, the drop in rate is much smaller in phase IV of campaign 3, with an exponential decay indicating the presence of radon remaining in the detector.
As an initial assessment of the trap performance, we compare the rates obtained in phases I and IV with either quality or 214Po cut throughout the three campaigns. The rate in phase I for all campaigns is approximately constant, and is fitted with a constant function. To compare this rate with the rate in phase IV after the trap is opened, we define the following reference time. We extrapolate the expected rate from phase III (without trap) with quality cut to later times and determine the reference time where the extrapolation reaches 76.5 Hz (dashed vertical lines in Fig. 2). This procedure is applied in all three campaigns to determine the rates in phases I and IV, see Table S1 of [SOM]. From this, we calculate the rate ratio between phase IV and phase I with either the quality or 214Po cut, as shown in Fig. 3.
Campaigns 1 and 2, using the silver-zeolite trap, consistently achieve near complete radon removal, whereas the radon levels in the activated charcoal campaign 3 are two to three orders of magnitude higher. As expected, the rate ratio obtained with the 214Po cut is typically larger, but more affected by statistical fluctuations due to the small rate of events in phase I and phase IV.
@@
Here, $R_{\text{exp}}(t)$ is the expected rate obtained from the extrapolated rate trend and $R_{\text{obs}}(t)$ is the observed rate from the experimental data after the trap is opened. The latter is modeled with a generic functional form over time, consisting of an exponential plus a constant rate, which is fit to the data using an MCMC [emcee].
The R-value at a given time in phase IV is determined by optimizing a Poisson likelihood (with a Nelder-Mead algorithm [Virtanen:2019joe, Gao:2012guu]) comparing the observed count model with the expected count rate scaled by $R(t)^{-1}$ [SOM]. The Feldman-Cousins method is applied to determine the 90% confidence level (CL) uncertainty bands for $R(t)^{-1}$ [Feldman:1997qc] over time. This process is repeated with many randomly–drawn MCMC samples for the extrapolated rate trend model as well as the observed rate model, and the resulting median value of the 90% CL intervals for $R(t)^{-1}$ is found for each time the calculation is performed (the central times of the bins used in Fig. 2).
Given the expected 15.1 cm average track length of 214Po events [Ziegler:2010] relative to the 30 cm diameter of the SPC, some events may not deposit all their energy within the detector volume. Further, some fraction of 214Po isotopes are expected to settle on the inner surface of the detector before decaying, increasing the fraction of events that will not deposit their full energy in the SPC gas. This could result in an underestimation of the actual rate of 214Po. Consequently, the expected and observed 214Po event rates in phase IV are both conservatively reduced by 50% in the calculation of $R(t)^{-1}$.
[…]
@@
In this Letter, we demonstrate that silver-zeolite Ag-ETS-10 achieves superior radon removal efficiency at room temperature compared to activated charcoal, which is typically used at cryogenic temperatures. Our innovative closed-loop system offers a novel approach beyond previous methods and is the first to demonstrate validated performance under realistic operating conditions, such as those present in dark-matter-type detectors. These findings not only confirm and extend previous studies [Obrien:2021, Heinitz:2023, 10.1093/ptep/ptad160, Ogawa:2022tku, Sone:2024gox, 10.1093/ptep/ptaf007], but also highlight the strong potential of silver-zeolite for ambient-temperature radon reduction systems in rare-event search experiments. This advancement is particularly important for underground laboratories, where elevated radon levels compromise the ultra-low background environments required for dark matter and neutrino physics experiments [Cui:2023fcd, Huang:2024]. Integrating silver-zeolite filters into ventilation systems and clean rooms can substantially reduce radon concentrations and enhance the supporting capabilities of these facilities [Kamaha:2022ooe, Baudis:2022qjb]. Moreover, eliminating the need for cryogenic cooling offers significant operational advantages by simplifying system design and reducing the overall complexity of radon reduction systems for future ton-scale detectors and beyond. The development of room-temperature radon filters could enable useful applications across multiple sectors, including residential basements, healthcare facilities, and other specialized industrial environments. Since our experimental campaigns with silver-zeolite Ag-ETS-10 have been conducted with an actual dark matter-type detector, they serve as a proof of principle for its application in current and future rare event searches.
Acknowledgments—This work was supported by the Canada First Research Excellence Fund through the Arthur B. McDonald Canadian Astroparticle Physics Research Institute and through the University of Alberta startup fund UOFAB Startup Piro. M.-C. P. also acknowledges funding from the Natural Sciences and Engineering Research Council of Canada (NSERC) Subatomic Physics Discovery Grant (project) No. SAPPJ-2024-00025, from which P. G. and Y. D. were further supported. The work of D. D. and C. G. was also supported by the NSERC Graduate Scholarships Doctoral and Master programs, respectively. We are grateful for the support of Compute Ontario [ontario], the BC DRI Group, and the Digital Research Alliance of Canada [alliance] required to carry out the simulation needed for this Letter. We thank Hervé Le Provost (IRFU-CEA/Saclay) for his software development work in our data acquisition system. We also thank the NEWS-G collaboration for their technical support provided for this work and the review of this Letter. We also thank Daniel Kuznicki for assisting with the silver-zeolite Ag-ETS-10 sample obtained from Extraordinary Adsorbents Inc. of Edmonton, Alberta [ExtraEdmonton1].
@@
This Supplemental Material provides further technical details regarding the regeneration process of the adsorbents, the cuts used in our data analysis, the expected decay rates derived with a Monte Carlo (MC) model and Markov Chain Monte Carlo (MCMC) fit, the rates in phases I and IV obtained throughout the campaigns, and calculation of the R-values.